# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 5 Outlook on IB-MOT and applications
In order to construct a FAM $(M_{t})$, one needs a martingale vector $X$ and an $X$-RAP. In some real-world situations, one might not have the vector $X$, but only its marginal distributions. In such a case, one must select a martingale coupling for these marginal distributions in order to construct a FAM. Could the behaviour of the FAMs associated with each possible coupling for the given marginal distributions inform us on which coupling to choose, i.e., is there an optimal coupling for a FAM? For each coupling, a FAM produces the best estimate of the last target random variable given the information generated by the $X$-RAP up until the time at which the estimation is performed. A natural question arises: Among all the estimates, which coupling yields the worst-case scenario? This idea of choosing a coupling based on paths connecting its marginal distributions on the space of probability measures reminds us of optimal transport. We flesh out our thinking by focusing on FAMs in the case of $n=1$.
@@
Optimal transport (OT) dates back to Gaspard Monge in 1781 [23], with significant advancements by Leonid Kantorovich in 1942 [13] and Yann Brenier in 1987 [6]. It provides a way of comparing two measures, $\mu$ and $\nu$, defined on the Borel sets of topological Hausdorff spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We denote by $\mathcal{P}(\mathcal{X})$ and $\mathcal{P}(\mathcal{Y})$ the sets of Borel measures on $\mathcal{X}$ and $\mathcal{Y}$ respectively. A popular mental image in the context of optimal transport is a pile of sand, modelled by a measure $\mu$, and a hole, modelled by another measure $\nu$. One wishes to fill the hole with the sand in an optimal manner, by exercising the least amount of effort. To make this statement more precise, one needs a cost function $c:\mathcal{X}\times\mathcal{Y}\rightarrow[0,\infty]$ that measures the cost of transporting a unit mass from $x\in\mathcal{X}$ to $y\in\mathcal{Y}$. The optimal transport problem is concerned with how to transport $\mu$ to $\nu$ whilst minimizing the cost of transportation. That is, given $\mu\in\mathcal{P}(\mathcal{X})$ and $\nu\in\mathcal{P}(\mathcal{Y})$,
$\inf_{\pi\in\Pi(\mu,\nu)}\textup{{K}}(\pi):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,\mathrm{d}\pi(x,y).$ (5.1)
This problem possesses many interesting properties. For instance, if $\mathcal{X}=\mathcal{Y}$ is a Polish space, and $(\mathcal{X},c)$ is a metric space, then
$W_{p}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{X}}c(x,y)^{p}\mathrm{d}\pi(x,y)\right)^{1/p}$ (5.2)
[…]
@@
is the shortest paths from $\mu$ to $\nu$ in $(\mathcal{P}_{p}(\mathcal{X}),W_{p})$. This links optimal transport to interpolation on the space of random variables. There are two main differences between $(M_{t})$ and $(X_{0}(T_{1}-t)/(T_{1}-T_{0})+X_{1}(t-T_{0})/(T_{1}-T_{0}))$. The latter process is not a martingale, and does not incorporate noise, resulting in a deterministic interpolation conditionally on $(X_{0},X_{1})$. Could we modify OT in a way to incorporate noise, resulting in a martingale interpolator instead? We treat these two differences separately.
To achieve interpolation in the presence of noise within the optimal transport framework, the entropic regularization of optimal transport provides a way forward. Given $\mu\in\mathcal{P}(\mathcal{X})$, $\nu\in\mathcal{P}(\mathcal{Y})$, and $\varepsilon>0$, one considers
$\inf_{\pi_{\varepsilon}\in\Pi(\mu,\nu)}\textup{{K}}_{\varepsilon}(\pi_{\varepsilon}):=\inf_{\pi_{\varepsilon}\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,\mathrm{d}\pi_{\varepsilon}(x,y)+\varepsilon\textup{ KL}(\pi_{\varepsilon}\,|\,\mu\otimes\nu)$ (5.4)
where KL is the Kullback-Leibler divergence
[…]
$F_{S}(\cdot)=\int_{\mathcal{X}}\int_{\mathcal{X}}F(x,y,\cdot)\,\mathrm{d}\pi_{S}^{*}(x,y),$ (5.6)
where $F(x,y,\cdot)$ is the distribution function of a Brownian bridge starting at time $T_{0}$ in value $x$ and ending at time $T_{1}$ in value $y$, is called a Schrödinger bridge. Such a process is the $X$-RAP
$I_{t}^{(1)}=B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}(B_{T_{0}}-X_{0})-\frac{t-T_{0}}{T_{1}-T_{0}}(B_{T_{1}}-X_{1})$ (5.7)
for $(X_{0},X_{1})\sim\pi_{S}^{*}$, and where $(B_{t})$ is a standard Brownian motion. Another characterization of such a process is the diffusion distributed according to $\mu$ at time $T_{0}$ and to $\nu$ at time $T_{1}$ with constant volatility $\sigma^{2}=\varepsilon/2(T_{1}-T_{0})$ and drift with the smallest possible $L^{2}(\Omega\times[T_{0},T_{1}])$-norm.
Next, we address how to create a martingale interpolator in the optimal transport setting. To this end, we adapt optimal transport to yield a martingale coupling: Given $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R})$ in convex order,
$\inf_{\pi\in\mathcal{M}(\mu,\nu)}\textup{{K}}(\pi)=\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int_{\mathbb{R}^{2}}c(x,y)\,\mathrm{d}\pi(x,y).$ (5.8)
We denote the minimizers of this problem by $\pi_{m}^{*}$. There are many differences between optimal transport and its martingale counterpart. For instance, the canonical cost function $c(x,y)=(x-y)^{2}$ for optimal transport cannot be used in the martingale context because, in this case, the objective function $\textup{{K}}(\pi)$ does not depend on $\pi$:
$\int_{\mathbb{R}^{2}}(x-y)^{2}\,\mathrm{d}\pi(x,y)$ $=\int_{\mathbb{R}^{2}}x^{2}+y^{2}-2xy\,\mathrm{d}\pi(x,y)$
$=\int_{\mathbb{R}^{2}}x^{2}\,\mathrm{d}\mu(x)+\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y)-2\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y)$
$=\int_{\mathbb{R}^{2}}x^{2}\,\mathrm{d}\mu(x)-\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y).$ (5.9)
Another difference is that one loses the geodesic interpretation since there is no counterpart to the Wasserstein distance in the martingale context.
@@
Hence, when attempting to impose the martingale condition in OT, the interpolator corresponding to the coupling is lost, whereas when trying to inject noise in OT, the problem becomes the opposite of a martingale problem: one seeks to find a drift for a diffusion with a constant volatility. It is our view that FAMs could be used to reconcile the introduction of noise in martingale optimal transport (MOT) and martingale interpolation. Instead of comparing the distributions of $X_{0}$ and $X_{1}$ using the expectation $\mathbb{E}_{\pi}[c(X_{0},X_{1})]$, we propose to investigate the expectation of the cumulative cost in time on $[T_{0},T_{1}]$ incurred by estimating $X_{1}$ by $M_{t}$. This would result in an information-based version of MOT, since the information flow generated by the $X$-RAP is taken into account in determining the optimal coupling. Because we consider the cumulative cost in time on $[T_{0},T_{1}]$, we propose to multiply the cost by a weight function that informs the problem about the rate of convergence of $M_{t}$ towards $X_{1}$. By construction, the cost will always be small towards the end of the time interval, regardless of the coupling. The weight function allows for a fair comparison between the costs arising across different times by removing the bias introduced in the construction of $(M_{t})$.
This information-based version of MOT may be useful in an asset pricing context, for instance. A term structure of options allows to imply the market-consistent marginal distributions of the underlying asset at the fixed option expiry dates. Information-based MOT provides a method to explicitly construct the optimal no-arbitrage dynamics in continuous-time of an asset price process, e.g., the price process of a stock, that interpolates between the option-implied marginal distributions, i.e., the optimal FAM. By maximizing the expectation of the cumulative cost in time between $X_{1}$ and $M_{t}$ over the set of martingale couplings, and constructing the FAM corresponding to the maximizing martingale coupling, one obtains the ”worst case” peacock given the information generated by the underlying RAP, allowing us to establish probabilistic price bounds for the stock. The theory on, and applications thereof, information-based martingale optimal transport (IB-MOT) is developed in a follow-on paper [15].
@@
Information-based stochastic interpolation can also be considered using FARMs $(M_{t}^{-})$. For example, a governmental body may need to decide on future climate goals in terms of, e.g., atmospheric CO2 emissions. Future goals depend on a succession of intermediate achievements, which are reached with uncertainty. So, these goals can be modelled by random variables with probability distributions in concave order, if a governmental body is interested in progressively decreasing the uncertainty around the mean until the final goal is reached. One can then look for an optimal reverse-martingale coupling linking the intermediate goals using the same procedure described above while replacing FAMs with FARMs: one aims at maximizing the expectation of the cumulative cost in time between $X_{0}$ and $M_{t}^{-}$ over the set of reverse-martingale couplings. Then, constructing the FARM corresponding to the maximizing reverse-martingale coupling, one obtains the ”worst case” reverse-peacock given the information generated by the underlying RAP, allowing one to understand if the goals are likely to be achieved, and if so, what trajectories must be taken in the near future to succeed.
@@ 6 Conclusions
Given a finite sequence of convexly ordered random variables that are indexed by pre-specified fixed dates, we show how continuous-time martingales, the so-called filtered arcade martingales (FAMs, Definitions 4.1 and 4.22) can be constructed explicitly, such that the target random variables are matched, at the given dates, almost surely. These interpolating martingales, defined as conditional expectations, filter noisy information about the last target random variable such that the interpolation through the whole sequence of random variables is guaranteed. This martingale-interpolating method relies on a specific class of stochastic processes, the randomized arcade processes (RAPs, Definition 3.1). These processes, which interpolate strongly between the target random variables, generate information flows that reveal each target random variable at its indexing date, thus allowing a FAM to interpolate through the sequence of random variables. FAMs are solutions to filtering problems where the best estimate of the last target random variable is calculated given the filtration generated by the observation process, i.e., the RAP.
@@
A randomized arcade process is the sum of two independent components: a signal function and a noise process, the latter is the so-called arcade process (AP, Definition 2.2). An AP is a functional of a stochastic driving process, and interpolates strongly between zeros. A well-known example of an AP over one period of time is the anticipative Brownian bridge (Example 2.3). Since APs play the role of noise, a natural class of APs are the ones driven by a Gaussian process. We give sufficient condition for such an AP to be Markovian (Corollary 2.11), and show that there are Markovian APs that are not anticipative representations of Markov bridges (Example 2.18). As for the signal function, it interpolates between the target random variables, and it is deterministic conditional on these random variables. Since a RAP is a sum of a signal function and an AP, it does not possess the Markov property with respect to its natural filtration, apart from in trivial cases. Instead, it may have a similar property, the conditionally-Markov property (Definition 3.7), which plays a central role for the computational tractability of a FAM, and for guaranteeing the interpolation in the case with more than two random variables. We give sufficient conditions for a RAP to satisfy the conditionally-Markov property (Theorem 3.8), and we consider FAMs constructed with such RAPs. Using the Bayes rule and Itô’s lemma, we derive the SDEs satisfied by such FAMs (Propositions 4.5 and 4.8). If furthermore, the RAPs are standard (Definition 3.6), the obtained FAMs can be expressed in terms of an integral with respect to a Brownian motion adapted to the filtration generated by the RAP underlying the FAM (Proposition 4.13). This Brownian motion corresponds to the innovations process in stochastic filtering theory.
@@
The FAM theory relies on convexly ordered probability measures, that is, the existence of a martingale coupling for the target variables. In some real life scenarios, one might not have the coupling, but only its marginal distributions. In that case, one must select a martingale coupling for these marginal distributions in order to construct a FAM. As a potential application in the field of optimal transport, that will be studied in a followup paper, we propose the use of the paths of FAMs to select an optimal coupling for the marginal distributions. This method is inspired by martingale optimal transport and Schrödinger’s problem. It aims at introducing noise in MOT in a similar fashion to how the entropic regularization of optimal transport introduces noise in optimal transport. The noise contained in FAMs, which depends on the choice of the underlying RAP, is taken into consideration in the selection of an optimal coupling for the target random variables: we aim at maximizing the expectation of the cumulative cost in time between the final target random variable and the considered FAM over the set of martingale couplings. The same ideas can be applied to FARMs, where we look instead for an optimal reverse-martingale coupling based on the paths of FARMs. Arcade processes and randomized arcade processes, alongside the filtered arcade martingales, appear thereof to suggest applications in, e.g., financial and insurance mathematics, mathematical biology, statistics, and climate science.

### v3

@@ 6 Outlook on IB-MOT and applications
In this paper, a new approach is developed that allows for strong interpolation of a sequence of random variables by deploying a generalisation of anticipative stochastic bridges, the RAPs. These processes generate filtrations that are used to construct FAMs, a class of processes that interpolate between the components of a discrete-time martingale, almost surely. For a fixed driver and interpolating coefficients, there is essentially a one-to-one mapping between discrete-time martingales and FAMs, i.e., no two FAMs with the same driver and interpolating coefficients interpolate the same discrete-time martingale.
FAMs solve the so-called martingale interpolating problem on the space of random variables in a novel way. Relaxing this interpolation problem on the space of probability measures provides a new degree of freedom for FAMs: in law, there are as many FAMs sharing the same driver, interpolating coefficients, and connecting the target measures, as there are elements in the set of martingale couplings of the target measures. Selecting one of the FAMs in law amounts to selecting one of the martingale couplings. This is reminiscent of optimal transport (see [25], [14], [6] for foundational work and [5], [33], [34] for modern treatments), where the aim is to select an optimal coupling. The question of whether there is a connection to optimal transport and, if so, what relation it might be, arises naturally at this stage.
Hence, in what follows, we investigate FAMs in law in the context of optimal transport and its noisy versions, since FAMs are inherently noisy. This then serves as a segue into considering martingale transport within the filtered arcade martingale paradigm, which we call information-based martingale optimal transport (IB-MOT), given the FAMs’ in-built informational (Bayesian) updating of their trajectories to target. The development of IB-MOT is presented in the working paper [16].
@@
The construction of a FAM $(M_{t})$ requires a martingale vector $X$ and an $X$-RAP. In real-world situations, one might not have the vector $X$, but only its marginal distributions. In such a case, one needs to select a martingale coupling for these marginal distributions in order to construct a FAM in law. Could the behaviour of the FAMs associated with each possible coupling for the given marginal distributions inform us on which coupling to choose, i.e., is there an optimal coupling for a FAM? For each coupling, a FAM produces the best estimate of the last target random variable given the information generated by the $X$-RAP up until the time at which the estimation is performed. A natural question arises: among all the estimates, which coupling yields the worst-case scenario? This idea of choosing a coupling based on paths connecting its marginal distributions on the space of probability measures reminds us of optimal transport-like problems. Let us briefly summarise, for the purpose of clarifying the context of our train of thought, the salient points of optimal transport.
@@
Optimal transport (OT) dates back to Gaspard Monge in 1781 [25], with significant advancements by Leonid Kantorovich in 1942 [14] and Yann Brenier in 1987 [6]. It provides a way of comparing two measures, $\mu$ and $\nu$, defined on the Borel sets of topological Hausdorff spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We denote by $\mathcal{P}(\mathcal{X})$ and $\mathcal{P}(\mathcal{Y})$ the sets of Borel measures on $\mathcal{X}$ and $\mathcal{Y}$, respectively. A popular mental image in the context of optimal transport is a pile of sand, modelled by a measure $\mu$, and a hole, modelled by another measure $\nu$. One wishes to fill the hole with the sand in an optimal manner, by exercising the least amount of effort. To make this statement more precise, one needs a cost function $c\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\times\mathcal{Y}\rightarrow[0,\infty]$ that measures the cost of transporting a unit mass from $x\in\mathcal{X}$ to $y\in\mathcal{Y}$. The optimal transport problem is concerned with how to transport $\mu$ to $\nu$ whilst minimising the cost of transportation. That is, given $\mu\in\mathcal{P}(\mathcal{X})$ and $\nu\in\mathcal{P}(\mathcal{Y})$,
$\inf_{\pi\in\Pi(\mu,\nu)}\textup{{K}}(\pi)\mathrel{\mathop{\ordinarycolon}}=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,\mathrm{d}\pi(x,y).$ (6.1)
This problem possesses many interesting properties. For instance, under mild conditions, it defines a collection of metrics denoted $W_{p}$ for any $p\geqslant 1$, the Wasserstein $p$-metric, on the space $\mathcal{P}_{p}(\mathcal{X})$ of probability measures on $\mathcal{X}$ with finite $p$th moment. Furthermore, if $\mathcal{X}$ is Euclidean, and $c(x,y)=\mathinner{\!\left\lVert x-y\right\rVert}$, there is a one-to-one correspondence between the minimisers $\pi^{*}$ of $\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)^{p}\,\mathrm{d}\pi(x,y)$, and the geodesics in $(\mathcal{P}_{p}(\mathcal{X}),W_{p})$. If $(X_{0},X_{1})\sim\pi^{*}$, then the law of the process
@@
is the shortest paths from $\mu$ to $\nu$ in $(\mathcal{P}_{p}(\mathcal{X}),W_{p})$. This links optimal transport to interpolation on the space of random variables. There are two main differences between a FAM for $(X_{0},X_{1})$ and $(X_{0}(T_{1}-t)/(T_{1}-T_{0})+X_{1}(t-T_{0})/(T_{1}-T_{0}))$. The latter process is not a martingale, and does not incorporate noise, resulting in a deterministic interpolation conditional on $(X_{0},X_{1})$. Could we modify OT in a way to incorporate noise and produce a martingale interpolator instead? Separately, these conditions are satisfied by well-studied offsprings of OT. The entropic regularisation of optimal transport ([19], [26]), which coincides with Schrödinger’s problem [28] under mild conditions, injects Brownian noise into OT, producing a noisy interpolator. On the other hand, martingale optimal transport ([2], [3]) replaces $\Pi(\mu,\nu)$ by $\mathcal{M}(\mu,\nu)$ in Eq. (6.1). However, when imposing the martingale condition in OT, the interpolator corresponding to the coupling is lost, whereas when trying to inject noise in OT, the problem becomes the opposite of a martingale problem: one seeks to find a drift for a diffusion with a constant volatility.
@@
It is our view that FAMs could be used to reconcile the introduction of noise in martingale optimal transport (MOT) and martingale interpolation. Instead of comparing the distributions of $X_{0}$ and $X_{1}$ using the expectation $\mathbb{E}_{\pi}[c(X_{0},X_{1})]$, we propose to investigate the expectation of the cumulative cost in time on $T_{0},T_{1}]$ incurred by estimating $X_{1}$ by $M_{t}$. This would result in an information-based version of MOT, since the information flow generated by the $X$-RAP is taken into account in determining the optimal coupling. Mathematically, the problem can be stated as
$\sup_{(X_{0},X_{1})\sim\pi^{X}\in\mathcal{M}(\mu,\nu)}\mathbb{E}\left[\int_{T_{0}}^{T_{1}}\frac{[X_{1}-M_{t}(X_{0},I_{t}^{(1)})]^{2}\sqrt{H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)}}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}\,\mathrm{d}t\right],$ (6.3)
see Section 5 for the introduced notion. Because we consider the cumulative cost in time on $T_{0},T_{1}]$, we propose to multiply the cost $X_{1}-M_{t}(X_{0},I_{t}^{(1)})]^{2}$ by a weight function that informs the problem about the rate of convergence of $M_{t}$ towards $X_{1}$. By construction, the cost will always be small towards the end of the time interval, regardless of the coupling. The weight function allows for a fair comparison between the costs arising across different times by removing the bias introduced in the construction of $(M_{t})$.
By maximising the expectation of the cumulative cost in time between $X_{1}$ and $M_{t}$ over the set of martingale couplings, and constructing the FAM corresponding to the maximising martingale coupling, one obtains the “worst-case” FAM in law given the information generated by the underlying RAP. The IB-MOT theory, and applications thereof, is developed in a follow-on paper [16].
@@
Information-based stochastic interpolation can also be considered using FARMs $(M_{t}^{-})$. For example, a government body may need to decide on future climate goals in terms of, e.g., atmospheric CO2 emissions. Future goals depend on a succession of intermediate achievements, which are reached with uncertainty. So, these goals can be modelled by random variables with probability distributions in concave order, if a governmental body is interested in progressively decreasing the uncertainty around the mean until the final goal is reached. One can then look for an optimal reverse-martingale coupling linking the intermediate goals using the same procedure described above while replacing FAMs with FARMs: one aims at maximising the expectation of the cumulative cost in time between $X_{0}$ and $M_{t}^{-}$ over the set of reverse-martingale couplings. Then, constructing the FARM corresponding to the maximising reverse-martingale coupling, one obtains the “worst-case” FARM given the information generated by the underlying RAP, allowing one to understand if the goals are likely to be achieved, and if so, what trajectories must be taken in the near future to succeed.
We note that IB-MOT is not the only optimal transport problem that incorporates noise while maintaining the martingale property and producing an interpolator. The martingale Benamou-Brenier problem (MBBP) [1] and Schrödinger’s volatility models (SVMs) [11] are examples of such problems. IB-MOT is an addition to this class of problems, relying on filtering to select the optimal coupling instead.
@@ 7 Conclusions
Given a finite sequence of convexly ordered random variables that are indexed by pre-specified fixed dates, we show how continuous-time martingales, the so-called filtered arcade martingales (FAMs, Definitions 5.27 and 5.44) can be constructed explicitly such that the target random variables are matched at the given dates, almost surely. These interpolating martingales, defined as conditional expectations, filter noisy information about the last target random variable such that the interpolation through the whole sequence of random variables is guaranteed. This martingale-interpolating method relies on a specific class of stochastic processes, the randomized arcade processes (RAPs, Definition 3.16). These processes, which interpolate strongly between the target random variables, generate information flows that reveal each target random variable at its indexing date, thus allowing a FAM to interpolate through the sequence of random variables. FAMs are solutions to filtering problems where the best estimate of the last target random variable is calculated given the filtration generated by the observation process, i.e., the RAP.
@@
A randomised arcade process is the sum of two independent components: a signal function and a noise process, the latter is the so-called arcade process (AP, Definition 2.2). An AP is a functional of a stochastic driving process, and interpolates strongly between zeros. A well-known example of an AP over one period of time is the anticipative Brownian bridge (Example 2.3). Since APs play the role of noise, a natural class of APs are the ones driven by a Gaussian process. We give sufficient condition for such an AP to be Markovian (Corollary 2.8), and show that there are Markovian APs that are not anticipative representations of Markov bridges (Example 2.15). As for the signal function, it interpolates between the target random variables, and it is deterministic conditional on these random variables. Since a RAP is a sum of a signal function and an AP, it does not possess the Markov property with respect to its natural filtration, except in trivial cases. Instead, it may have a similar property, the conditionally-Markov property (Definition 4.21), which plays a central role for the computational tractability of a FAM, and for guaranteeing the interpolation in the case of more than two random variables. We give sufficient conditions for a RAP to satisfy the conditionally-Markov property (Theorem 4.22), and we consider FAMs constructed with such RAPs. Using the Bayes rule and Itô’s lemma, we derive the SDEs satisfied by such FAMs (Propositions 5.31 and 5.32). If furthermore, the RAPs are standard (Definition 3.20), the obtained FAMs can be expressed in terms of an integral with respect to a Brownian motion adapted to the filtration generated by the RAP underlying the FAM (Proposition 5.37). This Brownian motion corresponds to the innovations process in stochastic filtering theory.
@@
The FAM theory relies on convexly ordered probability measures, that is, the existence of a martingale coupling for the target variables. In some real life scenarios, one might not have the coupling, but only its marginal distributions. In that case, one must select a martingale coupling for these marginal distributions in order to construct a FAM. As a potential application in the field of optimal transport, that is studied in [16], we propose the use of the paths of FAMs to select an optimal coupling for the marginal distributions. This method is inspired by martingale optimal transport and Schrödinger’s problem. It aims at introducing noise in MOT in a similar fashion to how the entropic regularisation of optimal transport introduces noise in optimal transport. The noise contained in FAMs, which depends on the choice of the underlying RAP, is taken into consideration in the selection of an optimal coupling for the target random variables—we aim at maximising the expectation of the cumulative cost in time between the final target random variable and the considered FAM over the set of martingale couplings. The same ideas can be applied to FARMs, where we look instead for an optimal reverse-martingale coupling based on the paths of FARMs. Arcade processes and randomised arcade processes, alongside the filtered arcade martingales, suggest applications in, e.g., financial and insurance mathematics, mathematical biology, statistics, and climate science.